OpenStudy (anonymous):

A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Part B How many miles are represented by 1/4 inch on the map?

OpenStudy (anonymous):

i thought for A 4.8 and for B 175

OpenStudy (anonymous):

am i correct?

OpenStudy (phi):

**How many inches on the map represent 60 miles? *** I would write a ratio \[ \frac{2 \ inches}{25 \ miles} = \frac{x \ inches}{60 \ miles} \]

OpenStudy (anonymous):

am i correct that its 4.8?

OpenStudy (phi):

multiply both sides by 60 to find x \[ \frac{2 \ inches}{25 \ miles} \cdot 60 \ miles = x\]

OpenStudy (phi):

simplify 2*60/25 to get 4.8

OpenStudy (anonymous):

thats 4.8

OpenStudy (phi):

so I would say 4.8 inches

OpenStudy (anonymous):

so then did i get part b correct as 175?

OpenStudy (phi):

we know 2 inches is 25 miles 1/4 inch is shorter than 2 inches, so the miles will be less than 25

OpenStudy (anonymous):

yes

OpenStudy (phi):

write the ratio \[ \frac{2 \ inches}{25 \ miles} = \frac{\frac{1}{4} \ inches}{x \ miles} \]

OpenStudy (phi):

to solve, you can first "flip" both sides \[ \frac{25 \ miles}{2 \ inches} = \frac{x \ miles}{\frac{1}{4} \ inches} \]

OpenStudy (phi):

now multiply both sides by 1/4

OpenStudy (anonymous):

no just working sorry im still here

OpenStudy (anonymous):

ok i did this but for the last answer i got 175 miles

OpenStudy (phi):

forget the ratios for the moment. if 2 inches is 25 miles what is 1 inch ?

OpenStudy (anonymous):

12.5

OpenStudy (phi):

if 1 inch is 12.5 miles what is 0.5 inches ?

OpenStudy (phi):

i.e what is 1/2 inch?

OpenStudy (anonymous):

6.25

OpenStudy (phi):

and 1/4 inch ?

OpenStudy (anonymous):

ummm 6?

OpenStudy (phi):

1 inch is 12.5 1/4 inch is ?

OpenStudy (anonymous):

is it 3.125??

OpenStudy (phi):

yes.

OpenStudy (phi):

using ratios \[ \frac{25 \ miles}{2 \ inches} = \frac{x \ miles}{\frac{1}{4} \ inches} \\ \frac{25 \ miles}{2 \ inches} \cdot \frac{1}{4} \ inches= \frac{x \ miles}{\frac{1}{4} \ inches} \cdot \frac{1}{4} \ inches \]

OpenStudy (phi):

you get 25/8 = x or x= 3.125

OpenStudy (anonymous):

ok thank you!!

OpenStudy (phi):

try to use common sense if you get stuck.

OpenStudy (anonymous):

i get it now thanks!